A block of mass M = 10 kg is sitting on a surface inclined at angle θ = 45°. Given that the
coefficient of static friction is μ = 0.5 between block and surface, what is the minimum
force F necessary to prevent slipping? What is the maximum force F that can be exerted
without causing the block to slip?
1
Expert's answer
2020-12-14T07:18:27-0500
"F=mg\sin{\theta}-\mu mg\cos{\theta}\\
F=(10)(9.8)(\sin{45}-0.5\cos{45})\\F=35\ N"